OpenStudy (anonymous):

Can anyone help me? What is the solution to the equation 3(2x+4) over 2 -9=3x?

OpenStudy (dumbcow):

\[\frac{3(2x+4)}{2} -9 = 3x\] is that correct?

OpenStudy (anonymous):

yup :)

OpenStudy (dumbcow):

well first thing i notice is that 2x+4 has a common factor of 2 2x +4 = 2(x+2) the outside 2 will then cancel with 2 in denominator leaving \[3(x+2) -9 = 3x\]

OpenStudy (anonymous):

so that's the answer?

OpenStudy (anonymous):

it's no solution?

OpenStudy (dumbcow):

um no i was just trying to help you simplify it you have to try to solve for x

OpenStudy (anonymous):

oh okay :D x=0 right?

OpenStudy (dumbcow):

nope...try plugging in 0 for x, you wont get same thing on both sides

OpenStudy (dumbcow):

3(x+2) = 3x + 6 combine like terms

OpenStudy (anonymous):

ohh so x=6?

OpenStudy (dumbcow):

you should notice that there is a 3x on both sides, this means that the x will cancel out if you subtract 3x from both sides leaving -3 = 0 this obviously is false so there is no solution

OpenStudy (anonymous):

Okay thanks!!!

OpenStudy (dumbcow):

if you think x=6, you don;t need to ask if it is right. just plug in 6 for x and test to see if you get same thing on both sides

OpenStudy (dumbcow):

you can always plug in a number for x to see if its a solution...this way you can know whether your answer is right or not